OpenStudy (anonymous):

Use synthetic division to divide F(x) = x^3+4x^2-9x-36 by (x-3) ??

jhonyy9 (jhonyy9):

x^3 +4x^2 -9x -36 / x-3 =x^2 +7x +12 x^3 -3x^2 ---------- 0 7x^2 -9x 7x^2 -21x ----------- 0 12x -36 12x -36 --------- 0 0

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

x^2(x+4)-9(x+4)=(x^2 -9)(x+4)=(x-3)(x+3)(x+4) so than this divided by (x+3) equal =(x-3)(x+4)

jhonyy9 (jhonyy9):

hope you will like ithese --- ...